# Average Winding Number for Determinantal Curves associated to 2-Matrix Models in the Class AIII

### v1

@@ 4.1 Asymptotic behaviour of the tails
as in (4.12). Hence, there exists constants $c_{1},c_{2}>0$ such that the terms decay exponentially as $\mathcal{O}(N^{c_{1}}e^{-c_{2}N^{2\varepsilon}})$.
@@ 4.2 Asymptotic expansion of the denominator
Consequently, there exists a constant $c_{n}>0$ so that
@@ 4.4 Asymptotic expansion of the sum at the edge
$\begin{split}&{\rm C}_{1}^{(N)}(p)\underset{N\to\infty}{=}N\frac{a^{\prime}(p)\overline{a(p)}|a(p)|^{2\gamma-2}+b^{\prime}(p)\overline{b(p)}|b(p)|^{2\gamma-2}}{|a(p)|^{2\gamma}+|b(p)|^{2\gamma}}\\ &+\left(\frac{a^{\prime}(p)}{a(p)}-\frac{b^{\prime}(p)}{b(p)}\right)\biggl[\frac{\gamma-2\delta-1}{2}\frac{|a(p)|^{2\gamma}-|b(p)|^{2\gamma}}{|a(p)|^{2\gamma}+|b(p)|^{2\gamma}}+\mathcal{O}\left(\frac{N^{6\varepsilon}}{\sqrt{N\tau_{0}(p)(1-\tau_{0}(p))}}\right)\biggl].\end{split}$ (4.63)
@@
${\rm C}_{1}^{(N)}(p_{0}e^{i\delta\varphi})+{\rm C}_{1}^{(N)}(p_{0}e^{-i\delta\varphi})$ (4.64)
@@
$\begin{split}&{\rm C}_{1}^{(N,0)}(p_{0}e^{i\delta\varphi})+{\rm C}_{1}^{(N,0)}(p_{0}e^{-i\delta\varphi})\\ +&\frac{imc_{1}-(m+1)c_{2}\delta\varphi}{ic_{1}\delta\varphi-c_{2}\delta\varphi^{2}}\biggl[\frac{\gamma-2\delta-1}{2}\frac{|c_{1}\delta\varphi^{m}|^{2\gamma}-|b(p_{0})+ib^{\prime}(p_{0})\delta\varphi|^{2\gamma}}{|c_{1}\delta\varphi^{m}|^{2\gamma}+|b(p_{0})+ib^{\prime}(p_{0})\delta\varphi|^{2\gamma}}+\mathcal{O}\left(\frac{N^{6\varepsilon}}{\sqrt{N|\delta\varphi|^{2m\gamma}}}\right)\biggl]\\ -&\frac{imc_{1}+(m+1)c_{2}\delta\varphi}{ic_{1}\delta\varphi+c_{2}\delta\varphi^{2}}\biggl[\frac{\gamma-2\delta-1}{2}\frac{|c_{1}\delta\varphi^{m}|^{2\gamma}-|b(p_{0})-ib^{\prime}(p_{0})\delta\varphi|^{2\gamma}}{|c_{1}\delta\varphi^{m}|^{2\gamma}+|b(p_{0})-ib^{\prime}(p_{0})\delta\varphi|^{2\gamma}}+\mathcal{O}\left(\frac{N^{6\varepsilon}}{\sqrt{N|\delta\varphi|^{2m\gamma}}}\right)\biggl]\\ =&{\rm C}_{1}^{(N,0)}(p_{0}e^{i\delta\varphi})+{\rm C}_{1}^{(N,0)}(p_{0}e^{-i\delta\varphi})-i\frac{(\gamma-2\delta-1)c_{2}}{2c_{1}}+\mathcal{O}\left(\frac{N^{6\varepsilon}}{\sqrt{N|\delta\varphi|^{2m\gamma}}}\right),\end{split}$ (4.65)
@@
${\rm C}_{1}^{(N,0)}(p)=N\frac{a^{\prime}(p)\overline{a(p)}|a(p)|^{2\gamma-2}+b^{\prime}(p)\overline{b(p)}|b(p)|^{2\gamma-2}}{|a(p)|^{2\gamma}+|b(p)|^{2\gamma}}.$ (4.66)
@@ 4.5 Discussion for small parameter amplitudes
$\begin{split}{\rm C}_{1}^{(N)}(p_{0}e^{i\delta\varphi})+{\rm C}_{1}^{(N)}(p_{0}e^{-i\delta\varphi})=&{\rm C}_{1}^{(N,0)}(p_{0}e^{i\delta\varphi})+{\rm C}_{1}^{(N,0)}(p_{0}e^{-i\delta\varphi})\\ &\hskip-56.9055pt+\left[\frac{a^{\prime}(p_{0}e^{i\delta\varphi})}{a(p_{0}e^{i\delta\varphi})}+\frac{a^{\prime}(p_{0}e^{-i\delta\varphi})}{a(p_{0}e^{-i\delta\varphi})}\right]\mathcal{O}(N^{\varepsilon/2+5\eta/4})+\mathcal{O}(\delta\varphi).\end{split}$ (4.80)
@@
$\begin{split}\mathbb{E}\left(\mathrm{Wind}_{N}\right)=&\left[\int_{{\rm S}^{1}_{\eta}}+\int_{{\rm S}^{1}\setminus{\rm S}^{1}_{\eta}}\right]{\rm C}_{1}^{(N)}(p)\mathrm{d}p\\ =&\int_{{\rm S}^{1}}{\rm C}_{1}^{(N,0)}(p)\mathrm{d}p+\frac{\gamma-2\delta-1}{2}\int_{{\rm S}^{1}_{\eta}}\left(\frac{a^{\prime}(p)}{a(p)}-\frac{b^{\prime}(p)}{b(p)}\right)\frac{|a(p)|^{2\gamma}-|b(p)|^{2\gamma}}{|a(p)|^{2\gamma}+|b(p)|^{2\gamma}}\mathrm{d}p\\ &+o(1).\end{split}$ (4.82)
@@ 5 Conclusion
In this work we have investigated two natural extensions of the Gaussian additive 2-matrix model describing a disordered one-dimensional Hamiltonian in the AIII symmetry class of the Altland-Zirnbauer classification.
First, we introduced a general complex Ginibre random field on the unit circle that recovers, as a special case, the standard Ginibre additive 2-matrix model while extending its analytical framework. To assess the universality of the Gaussian results, we, then, analysed a broader family of non-Gaussian models: Pólya additive 2-matrix models at finite matrix size, focusing on the Muttalib-Borodin subclass in the large-$N$ regime. Compared to the GinUE, these random matrix ensembles display different tail behaviour in their mean level density.
@@
which essentially corresponds to a Aharonov-Anandan phase angle. The subleading $\mathcal{O}\left(1\right)$ correction vanishes in the Gaussian case but generally persists for non-Gaussian Pólya ensembles, where it can be interpreted as a re-weighted winding number around the origin. Its magnitude depends on the tail behaviour and repulsion from the origin in the mean level density, both of which determine the typical distance of the determinantal curve from the origin.
A central open question is the universality of the winding-number distribution itself. Under mild regularity conditions, one may expect a Central Limit Theorem to hold, leading to an asymptotically Gaussian law as in the Ginibre additive 2-matrix model for which earlier derivations [10] shows that the standard deviation scales as $N^{1/4}$. However, the precise variance and subleading corrections are expected to depend on ensemble-specific details as already highlighted for the mean winding number. This layered asymptotic structure is reminiscent of that observed in large-$N$ free-energy expansions of Coulomb gases, where the leading order captures universal mean-field behaviour, while coefficients in higher-order terms are of geometric nature and encode increasingly model-dependent information, see for example [29, Th. 2].
Extending this analysis to other symmetry classes such as CII and BDI in one-dimension remains a challenging open direction. In these cases, the lack of real and quaternionic analogues of Pólya ensembles of multiplicative type, combined with the Pfaffian rather than determinantal structure of correlation functions, renders the analysis substantially more involved, see for example [30].
@@ Appendix A Complex integral identities
The evaluation of the three kernels (3.12-3.15) relies on a generalized version of the Andréief identity [31] derived in [28, Appendix C.1] and reviewed in [32]. For completeness we restate it here in a slightly more general form, allowing integration over an arbitrary measurable space.

### v2

@@ 4.1 Asymptotic behaviour of the tails
as in (4.12). Hence, there exist constants $c_{1},c_{2}>0$ such that the terms decay exponentially as $\mathcal{O}(N^{c_{1}}e^{-c_{2}N^{2\varepsilon}})$.
@@ 4.2 Asymptotic expansion of the denominator
Consequently, there exist constants $c_{n}>0$ so that
@@ 4.4 Asymptotic expansion of the sum at the edge
$\begin{split}&\mathrm{C}_{1}^{(N)}(p)\underset{N\to\infty}{=}N\frac{a^{\prime}(p)\overline{a(p)}|a(p)|^{2\gamma-2}+b^{\prime}(p)\overline{b(p)}|b(p)|^{2\gamma-2}}{|a(p)|^{2\gamma}+|b(p)|^{2\gamma}}\\ &+\left(\frac{a^{\prime}(p)}{a(p)}-\frac{b^{\prime}(p)}{b(p)}\right)\biggl[\frac{\gamma-2\delta-1}{2}\frac{|a(p)|^{2\gamma}-|b(p)|^{2\gamma}}{|a(p)|^{2\gamma}+|b(p)|^{2\gamma}}+\mathcal{O}\left(\frac{N^{6\varepsilon}}{\sqrt{N\tau_{0}(p)(1-\tau_{0}(p))}}\right)\biggl].\end{split}$ (4.63)
@@
$\mathrm{C}_{1}^{(N)}(p_{0}e^{i\delta\varphi})+\mathrm{C}_{1}^{(N)}(p_{0}e^{-i\delta\varphi})$ (4.64)
@@
$\begin{split}&\mathrm{C}_{1}^{(N,0)}(p_{0}e^{i\delta\varphi})+\mathrm{C}_{1}^{(N,0)}(p_{0}e^{-i\delta\varphi})\\ +&\frac{imc_{1}-(m+1)c_{2}\delta\varphi}{ic_{1}\delta\varphi-c_{2}\delta\varphi^{2}}\biggl[\frac{\gamma-2\delta-1}{2}\frac{|c_{1}\delta\varphi^{m}|^{2\gamma}-|b(p_{0})+ib^{\prime}(p_{0})\delta\varphi|^{2\gamma}}{|c_{1}\delta\varphi^{m}|^{2\gamma}+|b(p_{0})+ib^{\prime}(p_{0})\delta\varphi|^{2\gamma}}+\mathcal{O}\left(\frac{N^{6\varepsilon}}{\sqrt{N|\delta\varphi|^{2m\gamma}}}\right)\biggl]\\ -&\frac{imc_{1}+(m+1)c_{2}\delta\varphi}{ic_{1}\delta\varphi+c_{2}\delta\varphi^{2}}\biggl[\frac{\gamma-2\delta-1}{2}\frac{|c_{1}\delta\varphi^{m}|^{2\gamma}-|b(p_{0})-ib^{\prime}(p_{0})\delta\varphi|^{2\gamma}}{|c_{1}\delta\varphi^{m}|^{2\gamma}+|b(p_{0})-ib^{\prime}(p_{0})\delta\varphi|^{2\gamma}}+\mathcal{O}\left(\frac{N^{6\varepsilon}}{\sqrt{N|\delta\varphi|^{2m\gamma}}}\right)\biggl]\\ =&\mathrm{C}_{1}^{(N,0)}(p_{0}e^{i\delta\varphi})+\mathrm{C}_{1}^{(N,0)}(p_{0}e^{-i\delta\varphi})-i\frac{(\gamma-2\delta-1)c_{2}}{2c_{1}}+\mathcal{O}\left(\frac{N^{6\varepsilon}}{\sqrt{N|\delta\varphi|^{2m\gamma}}}\right),\end{split}$ (4.65)
@@
$\mathrm{C}_{1}^{(N,0)}(p)=N\frac{a^{\prime}(p)\overline{a(p)}|a(p)|^{2\gamma-2}+b^{\prime}(p)\overline{b(p)}|b(p)|^{2\gamma-2}}{|a(p)|^{2\gamma}+|b(p)|^{2\gamma}}.$ (4.66)
@@ 4.5 Discussion for small parameter amplitudes
$\begin{split}\mathrm{C}_{1}^{(N)}(p_{0}e^{i\delta\varphi})+\mathrm{C}_{1}^{(N)}(p_{0}e^{-i\delta\varphi})=&\mathrm{C}_{1}^{(N,0)}(p_{0}e^{i\delta\varphi})+\mathrm{C}_{1}^{(N,0)}(p_{0}e^{-i\delta\varphi})\\ &\hskip-56.9055pt+\left[\frac{a^{\prime}(p_{0}e^{i\delta\varphi})}{a(p_{0}e^{i\delta\varphi})}+\frac{a^{\prime}(p_{0}e^{-i\delta\varphi})}{a(p_{0}e^{-i\delta\varphi})}\right]\mathcal{O}(N^{\varepsilon/2+5\eta/4})+\mathcal{O}(\delta\varphi).\end{split}$ (4.80)
@@
$\begin{split}\mathbb{E}\left(\mathrm{Wind}_{N}\right)=&\left[\int_{{\rm S}^{1}_{\eta}}+\int_{{\rm S}^{1}\setminus{\rm S}^{1}_{\eta}}\right]\mathrm{C}_{1}^{(N)}(p)\mathrm{d}p\\ =&\int_{{\rm S}^{1}}\mathrm{C}_{1}^{(N,0)}(p)\mathrm{d}p+\frac{\gamma-2\delta-1}{2}\int_{{\rm S}^{1}_{\eta}}\left(\frac{a^{\prime}(p)}{a(p)}-\frac{b^{\prime}(p)}{b(p)}\right)\frac{|a(p)|^{2\gamma}-|b(p)|^{2\gamma}}{|a(p)|^{2\gamma}+|b(p)|^{2\gamma}}\mathrm{d}p\\ &+o(1).\end{split}$ (4.82)
@@ 5 Conclusion
In this work we have investigated two natural extensions of the Gaussian additive 2-matrix model describing a disordered one-dimensional Hamiltonian in symmetry class AIII of the Altland-Zirnbauer classification.
First, we introduced a general complex Ginibre random field on the unit circle that recovers, as a special case, the standard Ginibre additive 2-matrix model while extending its analytical framework. To assess the universality of the Gaussian results, we then analysed a broader family of non-Gaussian models: Pólya additive 2-matrix models at finite matrix size, focusing on the Muttalib-Borodin subclass in the large-$N$ regime. Compared to the GinUE, these random matrix ensembles exhibit different tail behaviour in their mean level density.
@@
which essentially corresponds to an Aharonov-Anandan geometric phase angle. The subleading $\mathcal{O}\left(1\right)$ correction vanishes in the Gaussian case but generally persists for non-Gaussian Pólya ensembles, where it can be interpreted as a re-weighted winding number around the origin. Its magnitude depends on the tail behaviour and repulsion from the origin in the mean level density, both of which determine the typical distance of the determinantal curve from the origin.
A central open question is the universality of the winding-number distribution itself. Under mild regularity conditions, one may expect a central limit theorem to hold, leading to an asymptotically Gaussian law as in the Ginibre additive 2-matrix model for which earlier derivations [hahnUniversalCorrelations2024] show that the standard deviation scales as $N^{1/4}$. However, the precise variance and subleading corrections are expected to depend on ensemble-specific details as already highlighted for the mean winding number. This layered asymptotic structure is reminiscent of that observed in large-$N$ free-energy expansions of Coulomb gases, where the leading order captures universal mean-field behaviour, while coefficients in higher-order terms are of geometric nature and encode increasingly model-dependent information, see for example [SerfatyGaussianFluctuationsCoulomb2023, Th. 2].
Extending this analysis to other symmetry classes such as CII and BDI in one dimension remains a challenging open direction. In these cases, the lack of real and quaternionic analogues of Pólya ensembles of multiplicative type, combined with the Pfaffian rather than determinantal structure of correlation functions, renders the analysis substantially more involved, see for example [hahnWindingNumberStatistics2023a].
@@ Appendix A Complex integral identities
The evaluation of the three kernels (3.12-3.15) relies on a generalized version of the Andréief identity [andreevNoteRelationEntre1886] derived in [kieburgDerivationDeterminantalStructures2010, Appendix C.1] and reviewed in [forresterMeetAndreiefBordeaux2019]. For completeness we restate it here in a slightly more general form, allowing integration over an arbitrary measurable space.